A uniform beam 6m long and weighing 4kg rest on support p and q placed left and right 1m from each end of the beam. Weight of mass 10kg and 8kg are placed near p and q respectively. On ecah end of the beam. Calculate the reation at p and q
Answer #1
125N, 95N
answered by: REINY PLEASE HELP ME!
Answer #2
R1+R2,=220 Take moment around Q Cwm=4*R1+80 Acwm=100*5+40*2=580 Cwm=Acwm 4R1+80= 580 4R1=500 R1=125 From R1+R2= 220 R2=95
answered by: Vick